\begin{thebibliography}{36}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Asghar et~al.(2018)Asghar, Poupart, Hoey, Jiang, and
  Mou}]{asghar2018affective}
Nabiha Asghar, Pascal Poupart, Jesse Hoey, Xin Jiang, and Lili Mou. 2018.
\newblock Affective neural response generation.
\newblock In \emph{European Conference on Information Retrieval}, pages
  154--166. Springer.

\bibitem[{Banerjee and Lavie(2005)}]{banerjee2005meteor}
Satanjeev Banerjee and Alon Lavie. 2005.
\newblock Meteor: An automatic metric for mt evaluation with improved
  correlation with human judgments.
\newblock In \emph{Proceedings of the acl workshop on intrinsic and extrinsic
  evaluation measures for machine translation and/or summarization}, pages
  65--72.

\bibitem[{Clark et~al.(2019)Clark, Khandelwal, Levy, and
  Manning}]{clark2019does}
Kevin Clark, Urvashi Khandelwal, Omer Levy, and Christopher~D Manning. 2019.
\newblock What does bert look at? an analysis of bert's attention.
\newblock \emph{arXiv preprint arXiv:1906.04341}.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: Pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Dziri et~al.(2018)Dziri, Kamalloo, Mathewson, and
  Zaiane}]{dziri2018augmenting}
Nouha Dziri, Ehsan Kamalloo, Kory~W Mathewson, and Osmar Zaiane. 2018.
\newblock Augmenting neural response generation with context-aware topical
  attention.
\newblock \emph{arXiv preprint arXiv:1811.01063}.

\bibitem[{Edunov et~al.(2019)Edunov, Baevski, and Auli}]{edunov-etal-2019-pre}
Sergey Edunov, Alexei Baevski, and Michael Auli. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1409} {Pre-trained language
  model representations for language generation}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4052--4059,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Ghosh et~al.(2017)Ghosh, Chollet, Laksana, Morency, and
  Scherer}]{ghosh-etal-2017-affect}
Sayan Ghosh, Mathieu Chollet, Eugene Laksana, Louis-Philippe Morency, and
  Stefan Scherer. 2017.
\newblock \href {https://doi.org/10.18653/v1/P17-1059} {Affect-{LM}: A neural
  language model for customizable affective text generation}.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 634--642,
  Vancouver, Canada. Association for Computational Linguistics.

\bibitem[{Holtzman et~al.(2019)Holtzman, Buys, Forbes, and
  Choi}]{holtzman2019curious}
Ari Holtzman, Jan Buys, Maxwell Forbes, and Yejin Choi. 2019.
\newblock The curious case of neural text degeneration.
\newblock \emph{arXiv preprint arXiv:1904.09751}.

\bibitem[{Huang et~al.(2018)Huang, Zaiane, Trabelsi, and
  Dziri}]{huang2018automatic}
Chenyang Huang, Osmar Zaiane, Amine Trabelsi, and Nouha Dziri. 2018.
\newblock Automatic dialogue generation with expressed emotions.
\newblock In \emph{Proceedings of the 2018 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 2 (Short Papers)}, pages 49--54.

\bibitem[{Ippolito et~al.(2019)Ippolito, Kriz, Sedoc, Kustikova, and
  Callison-Burch}]{ippolito-etal-2019-comparison}
Daphne Ippolito, Reno Kriz, Joao Sedoc, Maria Kustikova, and Chris
  Callison-Burch. 2019.
\newblock \href {https://www.aclweb.org/anthology/P19-1365} {Comparison of
  diverse decoding methods from conditional language models}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 3752--3762, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Kincaid et~al.(1975)Kincaid, Fishburne~Jr, Rogers, and
  Chissom}]{kincaid1975derivation}
J~Peter Kincaid, Robert~P Fishburne~Jr, Richard~L Rogers, and Brad~S Chissom.
  1975.
\newblock Derivation of new readability formulas (automated readability index,
  fog count and flesch reading ease formula) for navy enlisted personnel.

\bibitem[{Li et~al.(2016{\natexlab{a}})Li, Galley, Brockett, Gao, and
  Dolan}]{diversity2016}
Jiwei Li, Michel Galley, Chris Brockett, Jianfeng Gao, and Bill Dolan.
  2016{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/N16-1014} {A diversity-promoting
  objective function for neural conversation models}.
\newblock In \emph{Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 110--119. Association for Computational Linguistics.

\bibitem[{Li et~al.(2016{\natexlab{b}})Li, Galley, Brockett, Gao, and
  Dolan}]{li-etal-2016-diversity}
Jiwei Li, Michel Galley, Chris Brockett, Jianfeng Gao, and Bill Dolan.
  2016{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/N16-1014} {A diversity-promoting
  objective function for neural conversation models}.
\newblock In \emph{Proceedings of the 2016 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 110--119, San Diego, California. Association for
  Computational Linguistics.

\bibitem[{Li et~al.(2016{\natexlab{c}})Li, Galley, Brockett, Spithourakis, Gao,
  and Dolan}]{li-etal-2016-persona}
Jiwei Li, Michel Galley, Chris Brockett, Georgios Spithourakis, Jianfeng Gao,
  and Bill Dolan. 2016{\natexlab{c}}.
\newblock \href {https://doi.org/10.18653/v1/P16-1094} {A persona-based neural
  conversation model}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 994--1003,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Li et~al.(2016{\natexlab{d}})Li, Monroe, Ritter, Jurafsky, Galley,
  and Gao}]{Li-reinforce}
Jiwei Li, Will Monroe, Alan Ritter, Dan Jurafsky, Michel Galley, and Jianfeng
  Gao. 2016{\natexlab{d}}.
\newblock \href {https://doi.org/10.18653/v1/D16-1127} {Deep reinforcement
  learning for dialogue generation}.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pages 1192--1202. Association for Computational
  Linguistics.

\bibitem[{Lin(2004)}]{lin2004rouge}
Chin-Yew Lin. 2004.
\newblock Rouge: A package for automatic evaluation of summaries.
\newblock \emph{Text Summarization Branches Out}.

\bibitem[{Lipton et~al.(2018)Lipton, Li, Gao, Li, Ahmed, and
  Deng}]{lipton2018bbq}
Zachary Lipton, Xiujun Li, Jianfeng Gao, Lihong Li, Faisal Ahmed, and Li~Deng.
  2018.
\newblock Bbq-networks: Efficient exploration in deep reinforcement learning
  for task-oriented dialogue systems.
\newblock In \emph{Thirty-Second AAAI Conference on Artificial Intelligence}.

\bibitem[{Liu et~al.(2016)Liu, Lowe, Serban, Noseworthy, Charlin, and
  Pineau}]{dialogue-eval}
Chia-Wei Liu, Ryan Lowe, Iulian Serban, Mike Noseworthy, Laurent Charlin, and
  Joelle Pineau. 2016.
\newblock \href {https://doi.org/10.18653/v1/D16-1230} {How not to evaluate
  your dialogue system: An empirical study of unsupervised evaluation metrics
  for dialogue response generation}.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pages 2122--2132. Association for Computational
  Linguistics.

\bibitem[{Mei et~al.(2017)Mei, Bansal, and Walter}]{mei17}
Hongyuan Mei, Mohit Bansal, and Matthew~R. Walter. 2017.
\newblock Coherent dialogue with attention-based language models.
\newblock In \emph{Proceedings of the National Conference on Artificial
  Intelligence (AAAI)}, San Francisco, CA.

\bibitem[{Niu and Bansal(2018)}]{niu2018polite}
Tong Niu and Mohit Bansal. 2018.
\newblock Polite dialogue generation without parallel data.
\newblock \emph{Transactions of the Association of Computational Linguistics},
  6:373--389.

\bibitem[{Novikova et~al.(2017)Novikova, Du{\v{s}}ek, Cercas~Curry, and
  Rieser}]{novikova-etal-2017-need}
Jekaterina Novikova, Ond{\v{r}}ej Du{\v{s}}ek, Amanda Cercas~Curry, and Verena
  Rieser. 2017.
\newblock \href {https://doi.org/10.18653/v1/D17-1238} {Why we need new
  evaluation metrics for {NLG}}.
\newblock In \emph{Proceedings of the 2017 Conference on Empirical Methods in
  Natural Language Processing}, pages 2241--2252, Copenhagen, Denmark.
  Association for Computational Linguistics.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of the 40th annual meeting on association for
  computational linguistics}, pages 311--318. Association for Computational
  Linguistics.

\bibitem[{Pennebaker et~al.(2001)Pennebaker, Francis, and
  Booth}]{pennebaker2001linguistic}
James~W Pennebaker, Martha~E Francis, and Roger~J Booth. 2001.
\newblock Linguistic inquiry and word count: Liwc 2001.
\newblock \emph{Mahway: Lawrence Erlbaum Associates}, 71(2001):2001.

\bibitem[{Poria et~al.(2019)Poria, Majumder, Mihalcea, and
  Hovy}]{poria2019emotion}
Soujanya Poria, Navonil Majumder, Rada Mihalcea, and Eduard Hovy. 2019.
\newblock Emotion recognition in conversation: Research challenges, datasets,
  and recent advances.
\newblock \emph{arXiv preprint arXiv:1905.02947}.

\bibitem[{Radford et~al.(2019)Radford, Wu, Child, Luan, Amodei, and
  Sutskever}]{radford2019language}
Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, and Ilya
  Sutskever. 2019.
\newblock Language models are unsupervised multitask learners.
\newblock \emph{OpenAI Blog}, 1(8).

\bibitem[{Rashkin et~al.(2019)Rashkin, Smith, Li, and
  Boureau}]{rashkin2019towards}
Hannah Rashkin, Eric~Michael Smith, Margaret Li, and Y-Lan Boureau. 2019.
\newblock Towards empathetic open-domain conversation models: a new benchmark
  and dataset.
\newblock In \emph{ACL}.

\bibitem[{Santhanam and Shaikh(2019)}]{santhanam2019towards}
Sashank Santhanam and Samira Shaikh. 2019.
\newblock Towards best experiment design for evaluating dialogue system output.
\newblock \emph{arXiv preprint arXiv:1909.10122}.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch}]{sennrich-etal-2016-neural}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock \href {https://doi.org/10.18653/v1/P16-1162} {Neural machine
  translation of rare words with subword units}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  5998--6008.

\bibitem[{Venkatesh et~al.(2018)Venkatesh, Khatri, Ram, Guo, Gabriel, Nagar,
  Prasad, Cheng, Hedayatnia, Metallinou et~al.}]{venkatesh2018evaluating}
Anu Venkatesh, Chandra Khatri, Ashwin Ram, Fenfei Guo, Raefer Gabriel, Ashish
  Nagar, Rohit Prasad, Ming Cheng, Behnam Hedayatnia, Angeliki Metallinou,
  et~al. 2018.
\newblock On evaluating and comparing conversational agents.
\newblock \emph{arXiv preprint arXiv:1801.03625}.

\bibitem[{Vinyals and Le(2015)}]{vinyals2015neural}
Oriol Vinyals and Quoc Le. 2015.
\newblock A neural conversational model.
\newblock \emph{arXiv preprint arXiv:1506.05869}.

\bibitem[{Wolf et~al.(2019)Wolf, Sanh, Chaumond, and
  Delangue}]{wolf2019transfertransfo}
Thomas Wolf, Victor Sanh, Julien Chaumond, and Clement Delangue. 2019.
\newblock Transfertransfo: A transfer learning approach for neural network
  based conversational agents.
\newblock \emph{arXiv preprint arXiv:1901.08149}.

\bibitem[{Yang et~al.(2019)Yang, Dai, Yang, Carbonell, Salakhutdinov, and
  Le}]{yang2019xlnet}
Zhilin Yang, Zihang Dai, Yiming Yang, Jaime Carbonell, Ruslan Salakhutdinov,
  and Quoc~V Le. 2019.
\newblock Xlnet: Generalized autoregressive pretraining for language
  understanding.
\newblock \emph{arXiv preprint arXiv:1906.08237}.

\bibitem[{Zhang et~al.(2018)Zhang, Dinan, Urbanek, Szlam, Kiela, and
  Weston}]{zhang-etal-2018-personalizing}
Saizheng Zhang, Emily Dinan, Jack Urbanek, Arthur Szlam, Douwe Kiela, and Jason
  Weston. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-1205} {Personalizing dialogue
  agents: {I} have a dog, do you have pets too?}
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 2204--2213,
  Melbourne, Australia. Association for Computational Linguistics.

\bibitem[{Zhou et~al.(2018)Zhou, Huang, Zhang, Zhu, and
  Liu}]{zhou2018emotional}
Hao Zhou, Minlie Huang, Tianyang Zhang, Xiaoyan Zhu, and Bing Liu. 2018.
\newblock Emotional chatting machine: Emotional conversation generation with
  internal and external memory.
\newblock In \emph{Thirty-Second AAAI Conference on Artificial Intelligence}.

\bibitem[{Ziegler et~al.(2019)Ziegler, Melas-Kyriazi, Gehrmann, and
  Rush}]{ziegler2019encoder}
Zachary~M Ziegler, Luke Melas-Kyriazi, Sebastian Gehrmann, and Alexander~M
  Rush. 2019.
\newblock Encoder-agnostic adaptation for conditional language generation.
\newblock \emph{arXiv preprint arXiv:1908.06938}.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{acl2019}
\usepackage{times}
\usepackage{latexsym}
\usepackage{graphicx}
\usepackage{graphics}
\usepackage{url}
\usepackage{amsmath}
\usepackage{url}
\usepackage{enumitem}
\usepackage{color}
\usepackage{booktabs}
\usepackage{multirow}
\aclfinalcopy 

\usepackage{todonotes}
\usepackage{ftnxtra}

\title{Emotional Neural Language Generation Grounded in Situational Contexts}

\author{Sashank Santhanam  and Samira Shaikh\\
  Department of Computer Science \\
  University of North Carolina at Charlotte \\
  Charlotte, NC, USA\\
  \texttt{\{ssantha1,samirashaikh\}@uncc.edu} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
Emotional language generation is one of the keys to human-like artificial intelligence. Humans use different type of emotions depending on the situation of the conversation. Emotions also play an important role in mediating the engagement level with conversational partners. However, current conversational agents do not effectively account for emotional content in the language generation process. 
To address this problem, we develop a language modeling approach that generates affective content when the dialogue is situated in a given context. We use the recently released Empathetic-Dialogues corpus to build our models. Through detailed experiments, we find that our approach outperforms the state-of-the-art method on the perplexity metric by about 5 points and achieves a higher BLEU metric score. 
\end{abstract}

\section{Introduction}
Rapid advancement in the field of generative modeling through the use of neural networks has helped advance the creation of more intelligent conversational agents. Traditionally these conversational agents are built using \textit{seq2seq} framework that is widely used in the field of machine translation \cite{vinyals2015neural}. However, prior research has shown that engaging with these agents produces dull and generic responses whilst also being inconsistent with the emotional tone of conversation \cite{vinyals2015neural,li-etal-2016-persona}. These issues also affect engagement with the conversational agent, that leads to short conversations \cite{venkatesh2018evaluating}. Apart from producing engaging responses, understanding the situation and producing the right emotional response to a that situation is another desirable trait \cite{rashkin2019towards}. 

Emotions are intrinsic to humans and help in creation of a more engaging conversation \cite{poria2019emotion}. Recent work has focused on approaches towards incorporating emotion in conversational agents \cite{asghar2018affective,zhou2018emotional,huang2018automatic,ghosh-etal-2017-affect}, however these approaches are focused towards seq2seq task. We approach this problem of emotional generation as a form of transfer learning, using large pretrained language models. These language models, including BERT, GPT-2 and XL-Net, have helped achieve state of the art across several natural language understanding tasks \cite{devlin-etal-2019-bert,radford2019language,yang2019xlnet}. However, their success in language modeling tasks have been inconsistent \cite{ziegler2019encoder}. In our approach, we use these pretrained language models as the base model and perform transfer learning to fine-tune and condition these models on a given emotion. This helps towards producing more emotionally relevant responses for a given situation. In contrast, the work done by Rashkin \emph{et al.} \citeyearpar{rashkin2019towards} also uses large pretrained models but their approach is from the perspective of seq2seq task. 

Our work advances the field of conversational agents by applying the transfer learning approach towards generating emotionally relevant responses that is grounded on emotion and situational context. We find that our fine-tuning based approach outperforms the current state of the art approach on the automated metrics of the BLEU and perplexity. We also show that transfer learning approach helps produce well crafted responses on smaller dialogue corpus.

\section{Approach}
Consider the example show in Table \ref{problem_base} that shows a snippet of the conversation between a speaker and a listener that is grounded in a situation representing a type of emotion. Our goal is to produce responses to conversation that are emotionally appropriate to the situation and emotion portrayed.
\begin{table}[h]
\centering
\small
\begin{tabular}{@{}l@{}}
\toprule
\multicolumn{1}{l}{\textbf{Emotion: Confident}} \\ 
\multicolumn{1}{l}{\begin{tabular}[l]{@{}l@{}}\textbf{Situation:} I just knew I was going to do well at\\ work this morning.\end{tabular}} \\ \midrule
\begin{tabular}[c]{@{}l@{}}\textbf{Speaker:} I just knew I was going to do well at\\ work this morning. I was prepared\\ \textbf{Listener:} That is the way to go! Keep it up!\end{tabular} \\ \midrule
\end{tabular}
\caption{Example of conversations between a speaker and a listener}
\label{problem_base}
\end{table}
We approach this problem through a language modeling approach. We use large pre-trained language model 
as the base model for our response generation. This model is based on the transformer architecture and makes uses of the multi-headed self-attention mechanism to condition itself of the previously seen tokens to its left and produces a distribution over the target tokens. Our goal is to make the language model $p(y)=p(y_1,y_2,....,y_t;\theta)$ learn on new data and estimate the conditional probability $p(y|x)$. Radford \emph{et al.} 
\citeyearpar{radford2019language} demonstrated the effectiveness of language models to learn from a zero-shot approach in a multi-task setting. We take inspiration from this approach to condition our model on the task-specific variable $p(y_t|x,y_{< t})$, where $x$ is the task-specific variable, in this case the emotion label. We prepend the conditional variable (emotion, situational context) to the dialogue similar to the approach from Wolf \emph{et al} \citeyearpar{wolf2019transfertransfo}. We ensure that that the sequences are separated by special tokens.

\section{Experiments}
\subsection{Data}
In our experiments we use the Empathetic Dialogues dataset made available by Rashkin \emph{et al.} \citeyearpar{rashkin2019towards}. Empathetic dialogues is crowdsourced dataset that contains dialogue grounded in a emotional situation. The dataset comprises of 32  emotion labels including \textit{surprised, excited, angry, proud, grateful}. The speaker initiates the conversation using the grounded emotional situation and the listener responds in an appropriate manner\footnote{More information about the dataset made available on the \cite{rashkin2019towards}}.Table \ref{datasetstats} provides the basic statistics of the corpus. 
\begin{table}[htp]
\centering
\small
\begin{tabular}{cccc}
\toprule
\textbf{}           & \textbf{Train} & \textbf{Valid.} & \textbf{Test} \\
\toprule
\textbf{Num. Conversations}  & 19433           & 2770           & 2547          \\ \midrule
\textbf{Utterances} & 84324         & 12078           & 10973         \\ \midrule
\begin{tabular}[c]{@{}c@{}}\textbf{Avg Length}\\ \textbf{Conversations}\end{tabular}   & 4.31            & 4.36            & 4.31           \\ \midrule
\end{tabular}
\caption{Statistics of Empathetic Dialogue dataset used in our experiments}
\label{datasetstats}
\end{table}
\subsection{Implementation}
In all our experiments, we use the GPT-2 pretrained language model. We use the publicly available model containing 117M parameters with 12 layers; each layer has 12 heads. We implemented our models using PyTorch Transformers.\footnote{\url{https://github.com/huggingface/pytorch-transformers}} The input sentences are tokenized using byte-pair encoding(BPE) \cite{sennrich-etal-2016-neural} (vocabulary size of 50263). While decoding, we use the nucleus sampling ($p=0.9$) approach instead of beam-search to overcome the drawbacks of beam search \cite{holtzman2019curious,ippolito-etal-2019-comparison}. All our models are trained on a single TitanV GPU and takes around 2 hours to fine-tune the model. The fine-tuned models along with the configuration files and the code will be made available at: \url{https://github.com/sashank06/CCNLG-emotion}.

\subsection{Metrics}
Evaluating the quality of responses in open domain situations where the goal is not defined is an important area of research. Researchers have used methods such as BLEU , METEOR \cite{banerjee2005meteor}, ROUGE \cite{lin2004rouge} from machine translation and text summarization \cite{dialogue-eval} tasks. BLEU and METEOR are based on word overlap between the proposed and ground truth responses; they do not adequately account for the diversity of responses that are possible for a given input utterance and show little to no correlation with human judgments \cite{dialogue-eval}. We report on the BLEU \cite{papineni2002bleu} and Perplexity (PPL) metric to provide a comparison with the current state-of-the-art methods. We also report our performance using other metrics such as length of responses produced by the model. Following, Mei \emph{et al} \citeyearpar{mei17}, we also report the diversity metric that helps us measure the ability of the model to promote diversity in responses \cite{diversity2016}. Diversity is calculated as the as the number of distinct unigrams in the generation scaled by the total number of generated tokens \cite{mei17,li-etal-2016-persona}. We report on two additional automated metrics of readability and coherence. Readability quantifies the linguistic quality of text and the difficulty of the reader in understanding the text \cite{novikova-etal-2017-need}. We measure readability through the Flesch Reading Ease (FRE) \cite{kincaid1975derivation} which computes the number of words, syllables and sentences in the text. Higher readability scores indicate that utterance is easier to read and comprehend. Similarly, coherence measures the ability of the dialogue system to produce responses consistent with the topic of conversation. To calculate coherence, we use the method proposed by Dziri \emph{et al.} \citeyearpar{dziri2018augmenting}.

\section{Results}
\subsection{Automated Metrics}
\begin{table*}[h]
\centering
\small
\begin{tabular}{ccccccc}
\toprule
\textbf{Experiment}                                              & \textbf{\begin{tabular}[c]{@{}c@{}}Valid\\ PPL\end{tabular}} & \textbf{BLEU} & \textbf{Readability} & \textbf{Coherence} & \textbf{Length} & \textbf{Diversity} \\ \toprule
\begin{tabular}[c]{@{}c@{}}Baseline\\ Fine-Tuned \\ \cite{rashkin2019towards} \end{tabular}    & 21.24                                                        & 6.27          & x                    & x                  & x               & x                  \\ \midrule
\begin{tabular}[c]{@{}c@{}}Baseline\\ Emo-prepend \\ \cite{rashkin2019towards}\end{tabular}  & 24.30                                                        & 4.36          & x                    & x                  & x               & x                  \\ \midrule
\begin{tabular}[c]{@{}c@{}}Our Model\\ Fine-Tuned\end{tabular}   & \textbf{18.32}                                                        & 7.71          & 0.78                 & 0.93               & 9.77            & 0.0031             \\ \midrule
\begin{tabular}[c]{@{}c@{}}Our Model\\ Emo-prepend\end{tabular} & 19.49                                                        & \textbf{7.78}          & \textbf{0.79}                 & 0.93               & 9.71            & \textbf{0.0033}      \\ \bottomrule      
\end{tabular}
\caption{Comparison of the performance of our model to the baseline model proposed by Rashkin \emph{et al} \citeyearpar{rashkin2019towards} across a variety of automated metrics to provide a thorough comparison. \textbf{x} indicates that these metrics were not provided in the Rashkin \emph{et al} \citeyearpar{rashkin2019towards} work. }
\label{result_automated}
\end{table*}
We first compare the performance of our approach with the baseline results obtained from Rashkin \emph{et al.} \citeyearpar{rashkin2019towards} that uses a full transformer architecture \cite{vaswani2017attention}, consisting of an encoder and decoder. Table \ref{result_automated} provides a comparison of our approach with to the baseline approach. In Table \ref{result_automated}, we refer our ``\textit{Our Model Fine-Tuned}'' as the baseline fine-tuned GPT-2 model trained on the dialogue and ``\textit{Our-model Emo-prepend}'' as the GPT-2 model that is fine-tuned on the dialogues but also conditioned on the emotion displayed in the conversation. We find that fine-tuning the GPT-2 language model using a transfer learning approach helps us achieve a lower perplexity and a higher BLEU scores. The results from our approach are consistent with the empirical study conducted by Edunov \emph{et al} \citeyearpar{edunov-etal-2019-pre} that demonstrate the effectiveness of the using pre-trained model diminishes when added to the decoder network in an \textit{seq2seq} approach. We also perform a comparison between our two models on the metrics of length, diversity, readability and coherence. We find that our baseline model produces less diverse responses compared to when the model is conditioned on emotion. We find that the our \textit{emo-prepend} model also higher a slightly higher readability score that our baseline model.

\subsection{Qualitative Evaluation}
To assess the quality of generations, we conducted a MTurk human evaluation. We recruited a total of 15 participants and each participant was asked to evaluate 25 randomly sampled outputs from the test set on three metrics:
\begin{enumerate}[noitemsep,nolistsep]
    \item Readability - Is the response easy to understand, fluent and grammatical and does not have any consecutive repeating words.
    \item Coherence - Is the response relevant to the context of the conversation.
    \item Emotional Appropriateness- Does the response convey emotion suitable to the context of the conversation?
\end{enumerate}
\begin{table*}[h]
\centering
\small
\begin{tabular}{lll}
\toprule
\multicolumn{1}{c}{\textbf{\begin{tabular}[c]{@{}c@{}}Input \\ Conversation\end{tabular}}}                                                           & \multicolumn{1}{c}{\textbf{Model}}                     & \multicolumn{1}{c}{\textbf{Outputs}}                                                                     \\ \toprule
\multirow{3}{*}{\begin{tabular}[c]{@{}l@{}}Person A: I think I aced my  supervisor \\ exam!\end{tabular}}                                             & \begin{tabular}[c]{@{}l@{}}Ground\\ Truth\end{tabular} & \begin{tabular}[c]{@{}l@{}}Woo hoo! That is awesome! Congrats! When will you find out\end{tabular} \\ \cmidrule(l){2-3}
                                                                                                                                                     & \begin{tabular}[c]{@{}l@{}}Fine\\ Tuned\end{tabular}   & What was it?                                                                                             \\\cmidrule(l){2-3}
                                                                                                                                                     & \begin{tabular}[c]{@{}l@{}}Emo\\ Prepend\end{tabular}  & \begin{tabular}[c]{@{}l@{}}Wow, that's  impressive!\end{tabular}                               \\ \midrule
\multirow{3}{*}{\begin{tabular}[c]{@{}l@{}}Person A: I had pizza at a restaurant a\\ couple days ago and found  a hair on my \\pizza!\end{tabular}} & \begin{tabular}[c]{@{}l@{}}Ground\\ Truth\end{tabular} & \begin{tabular}[c]{@{}l@{}}It is disgusting. What did  you do?\end{tabular}                            \\ \cmidrule(l){2-3}
                                                                                                                                                     & \begin{tabular}[c]{@{}l@{}}Fine\\ Tuned\end{tabular}   & I was so angry                                                                                           \\ \cmidrule(l){2-3}
                                                                                                                                                     & \begin{tabular}[c]{@{}l@{}}Emo\\ Prepend\end{tabular}  & \begin{tabular}[c]{@{}l@{}}That's terrible,  did you clean it up?\end{tabular}        \\ \bottomrule         
\end{tabular}
\label{samples}
\caption{Example generations from our two model along with the ground truth responses. }
\end{table*}
\begin{table}[h]
\centering
\small
\begin{tabular}{p{1.6cm}p{1.4cm}p{1.4cm}p{1.5cm}}
\toprule
\textbf{}           & \textbf{Readability} & \textbf{Coherence} & \begin{tabular}[c]{@{}c@{}}\textbf{Emotional} \\ \textbf{Appropriate}-\\ \textbf{ness} \end{tabular} \\
\toprule
\begin{tabular}[c]{@{}c@{}}Our Model\\ Fine-Tuned\end{tabular} & 4.14           & 3.50           & 3.70          \\ \midrule
\begin{tabular}[c]{@{}c@{}}Our Model\\ Emo-prepend\end{tabular} & 3.54         & 3.4           & 3.19         \\ \midrule
\begin{tabular}[c]{@{}c@{}}Ground 
\\Truth \end{tabular}   & 3.92           & 3.86            & 4           \\ \midrule
\end{tabular}
\caption{Human ratings demonstrating a comparison between our models to the ground truth responses on the metrics of readability, coherence and emotional appropriateness}
\label{humaneval}
\end{table}
Table \ref{humaneval} shows the results obtained from the human evaluation comparing the performance of our fine-tuned, emotion pre-pend model to the ground-truth response. We find that our fine-tuned model outperforms the emo-prepend on all three metrics from the ratings provided by the human ratings.


\section{Related Work}
The area of dialogue systems has been studied extensively in both open-domain \cite{niu2018polite} and goal-oriented \cite{lipton2018bbq} situations. Extant approaches towards building dialogue systems has been done predominantly through the \textit{seq2seq} framework \cite{vinyals2015neural}. However, prior research has shown that these systems are prone to producing dull and generic responses that causes engagement with the human to be affected \cite{vinyals2015neural,venkatesh2018evaluating}. Researchers have tackled this problem of dull and generic responses through different optimization function such as MMI \cite{li-etal-2016-diversity} and through reinforcement learning approaches\cite{Li-reinforce}. Alternative approaches towards generating more engaging responses is by grounding them in personality of the speakers that enables in creating more personalized and consistent responses \cite{li-etal-2016-persona,zhang-etal-2018-personalizing,wolf2019transfertransfo}.

Several other works have focused on creating more engaging responses by producing affective responses. One of the earlier works to incorporate affect through language modeling is the work done by Ghosh 
\emph{et al.} \cite{ghosh-etal-2017-affect}. This work leverages the LIWC \cite{pennebaker2001linguistic} text analysis platform for affective features. Alternative approaches of inducing emotion in generated responses from a \textit{seq2seq} framework include the work done by Zhou \emph{et al}\citeyearpar{zhou2018emotional} that uses internal and external memory, Asghar \emph{et al.} \citeyearpar{asghar2018affective} that models emotion through affective embeddings and Huang 
\emph{et al} \citeyearpar{huang2018automatic} that induce emotion through concatenation with input sequence. More recently, introduction of transformer based approaches have helped advance the state of art across several natural language understanding tasks \cite{vaswani2017attention}. These transformers models have also helped created large pre-trained language models such as BERT \cite{devlin-etal-2019-bert}, XL-NET \cite{yang2019xlnet}, GPT-2 \cite{radford2019language}. However, these pre-trained models show inconsistent behavior towards language generation \cite{ziegler2019encoder}. 
\section{Conclusion and Discussion}
In this work, we study how pre-trained language models can be adopted for conditional language generation on smaller datasets. Specifically, we look at conditioning the pre-trained model on the emotion of the situation produce more affective responses that are appropriate for a particular situation. We notice that our fine-tuned and emo-prepend models outperform the current state of the art approach relative to the automated metrics such as BLEU and perplexity on the validation set. We also notice that the emo-prepend approach does not out perform a simple fine tuning approach on the dataset. We plan to investigate the cause of this in future work from the perspective of better experiment design for evaluation \cite{santhanam2019towards} and analyzing the models focus when emotion is prepended to the sequence \cite{clark2019does}. Along with this, we also notice other drawbacks in our work such as not having an emotional classifier to predict the outcome of the generated sentence, which we plan to address in future work. 
\label{conclusion}

 \section*{Acknowledgments}
This work was supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No FA8650-18-C-7881. All statements of fact, opinion or conclusions contained herein are those of the authors and should not be construed as representing the official views or policies of AFRL, DARPA, or the U.S. Government. We thank the anonymous reviewers for the helpful feedback. 

\bibliography{acl2019}
\bibliographystyle{acl_natbib}

\end{document}